OpenStudy (justin234):

What is the difference? 1/6 - 5/8 A. -19/24 B. -11/24 C. 11/24 D. 2

OpenStudy (tkhunny):

Have you considered finding a common denominator and completing the subtraction?

OpenStudy (mrnood):

Since the answers mainly have a denominator of 24 you might start with that so you need to convert the 2 terms into terms with 24 on the bottom.. \[\frac{ 1 }{ 6 } = \frac{ ? }{ 24 }\] \[\frac{ 5 }{ 8 } = \frac{ ? }{ 24 }\]

OpenStudy (mrnood):

@Justin234 you need to reply and contribute to get further help...